OpenStudy (anonymous):

A box contains 100 colored chips; some are yellow and some are green. John chooses a chip at random, records the color, and places it back in the bag. John has recorded 57 yellow chips and 19 green chips. Using these results, what is the predicted number of green chips in the box? 25 43 76 81

OpenStudy (anonymous):

anyone?

OpenStudy (lalaland_lauren):

What do you think it is?

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

can i give me the answer please iv been stuck on it for 10 minutes now

OpenStudy (lalaland_lauren):

I can't tell the answer for sure, but think about it. Because John pulled out more yellow than green, then there's probably only a few green chips in the bag.

OpenStudy (anonymous):

wait is it 25

OpenStudy (anonymous):

he picked 19 + 57 = 76 in total. 19 of them were green. This ratio will be the same for 100 chips in the bag. \[\frac{ 19 }{ 76 }=\frac{ x }{ 100 }\]

OpenStudy (anonymous):

yes 25

OpenStudy (lalaland_lauren):

Agree

OpenStudy (anonymous):

thank you! can you help me with a few others?

OpenStudy (lalaland_lauren):

Sure, go ahead

OpenStudy (anonymous):

Henry tossed a coin several times and recorded the number of heads and tails in the following table: side of coin: heads tails number of tosses for heads: 12 number of tosses for tails: 18 What is the experimental probability of getting heads? 0.4 0.7 1.2 1.8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

hello??

OpenStudy (lalaland_lauren):

sorry, something happened with my internet and everything went down

OpenStudy (lalaland_lauren):

But basically, since there were 30 flips of the coin in all, you have to divide the number of tries with heads vs total number of tries, like peachpi said earlier

OpenStudy (lalaland_lauren):

so your answer ends up being 0.4